5.For \\(g(x)=\\frac{x-4}{x-3}, we can use the mean value theorem on [4, 6], Hence determine \\(c\\)
a.\\(\\sqrt (112) \\)
b.\\(c=2\\pm \\sqrt(3)}\\)
c.\\(c=-2\\pm \\sqrt(5)}\\)
d.\\(c=3\\pm \\sqrt(3)}\\)

6.Given\\(f(x)=\\sqrt(9-x^{2})\\)
a.\\(\\frac{-3}{8}\\)
b.\\(\\frac{5}{8}\\)
c.\\(\\frac{7}{8}\\)
d.\\(\\frac{-9}{8}\\)
1
Expert's answer
2019-09-18T11:47:31-0400

QUESTION 5


Since the condition of the question says that we can use the mean value theorem on "[4,6]", we will not prove it, but only use the formula


"f'(c)=\frac{f(b)-f(a)}{b-a}"

( More information: https://en.wikipedia.org/wiki/Mean_value_theorem )


In our case,


"f(x)=\frac{x-4}{x-3}\rightarrow\begin{cases}
f(6)=\displaystyle\frac{6-4}{6-3}=\frac{2}{3}\\[0.5cm]
f(4)=\displaystyle\frac{4-4}{4-3}=0\\[0.5cm]
f'(x)=\displaystyle\frac{(x-4)'(x-3)-(x-4)(x-3)'}{(x-3)^2}=\frac{1}{(x-3)^2}
\end{cases}"

Then,



"f'(c)=\frac{f(6)-f(4)}{6-4}\rightarrow
\frac{1}{(c-3)^2}=\frac{\displaystyle\frac{2}{3}-0}{2}\rightarrow (c-3)^2=3\\[0.5cm]
|c-3|=\sqrt{3}\rightarrow c=3+\sqrt{3}\,\,\,{or}\,\,\,c=3-\sqrt{3}"

Since "c\in[4,6]" , then


"\boxed{c=3+\sqrt{3}}"

ANSWER

"\boxed{d.\,\,\,c=3\pm\sqrt{3}}"

Hint: We choose this answer, since only it contains the correct answer, although it also has an extra root.


QUESTION 6


Since the condition of the question is only "f(x)=\sqrt{9-x^2}" , and there is no continuation, then I cannot answer this question.